Solve using substitution
3x+y=13
2x-4y=18

Answer by rfer(16322)

y=13-3x
2x-4(13-3x)=18
2x-52+12x=18
14x=70
x=5
-----------
2*5-4y=18
10-4y=18
-4y=8
y=-2
-----------
3(5)-2=13